# S, T, U Parameters in The B-LSSM

### v1

@@ Appendix B Slef-Energys of Charginos and Neutralinos
$+(A_{L}B_{R}+A_{R}B_{L})\int\frac{d^{D}k}{(2\pi)^{D}}\frac{2m_{1}m_{2}}{d_{0}d_{1}}g^{\mu\nu}$
Where, $A^{\mu}=A_{L}P_{L}\gamma^{\mu}+A_{R}P_{R}\gamma^{\mu}+$. We define
@@
$=$ $i\frac{1-\ln(2\pi)\epsilon}{16\pi^{2}}\mathcal{C}_{2}2m_{1}m_{2}B_{0}(q^{2})g^{\mu\nu}$
$\mathcal{C}_{1}=$ $(A_{L}B_{L}+A_{R}B_{R})$
$\mathcal{C}_{2}=$ $(A_{L}B_{R}+A_{R}B_{L})$ (105)
@@
$\mathcal{M}_{1}(q^{2}=0)\sim$ $\frac{i}{16\pi}\mathcal{C}_{1}(m_{1}^{2}+m_{2}^{2})g^{\mu\nu}$
$\mathcal{M}_{2}(q^{2}=0)\sim$ $2\frac{i}{16\pi}\mathcal{C}_{2}m_{1}m_{2}g^{\mu\nu}$
$\frac{d\mathcal{M}(q^{2}=0)}{q^{2}}\sim$ $\frac{i}{16\pi}\frac{2}{3}\mathcal{C}_{1}g^{\mu\nu}$ (106)
@@
$\widetilde{\mathcal{C}}_{1\chi^{0}}^{ZZ}=$ $-2\frac{G^{2}}{4}(N^{*}_{j3}N_{i3}-N^{*}_{j4}N_{i4})^{2}$
$\widetilde{\mathcal{C}}_{2\chi^{0}}^{ZZ}=$ $2\frac{G^{2}}{4}(N^{*}_{j3}N_{i3}-N^{*}_{j4}N_{i4})^{2}$

### v2

@@ Appendix B SELF-ENERGIES of Charginos and Neutralinos
$+(A_{L}B_{R}+A_{R}B_{L})\int\frac{d^{D}k}{(2\pi)^{D}}\frac{2m_{1}m_{2}}{d_{0}d_{1}}g^{\mu\nu}.$
Where, $A^{\mu}=A_{L}P_{L}\gamma^{\mu}+A_{R}P_{R}\gamma^{\mu}$. We define
@@
$=$ $i\frac{1-\ln(2\pi)\epsilon}{16\pi^{2}}\mathcal{C}_{2}2m_{1}m_{2}B_{0}(q^{2})g^{\mu\nu},$
$\mathcal{C}_{1}=$ $(A_{L}B_{L}+A_{R}B_{R}),$
$\mathcal{C}_{2}=$ $(A_{L}B_{R}+A_{R}B_{L}).$ (54)
@@
$\mathcal{M}_{1}(q^{2}=0)\sim$ $\frac{i}{16\pi}\mathcal{C}_{1}(m_{1}^{2}+m_{2}^{2})g^{\mu\nu},$
$\mathcal{M}_{2}(q^{2}=0)\sim$ $2\frac{i}{16\pi}\mathcal{C}_{2}m_{1}m_{2}g^{\mu\nu},$
$\frac{d\mathcal{M}(q^{2}=0)}{q^{2}}\sim$ $\frac{i}{16\pi}\frac{2}{3}\mathcal{C}_{1}g^{\mu\nu}.$ (55)
@@
$+4c_{W}^{4}V^{*}_{j1}V_{i1}U_{j1}U^{*}_{i2}+c_{2W}^{2}V^{*}_{j2}V_{i2}U_{j2}U^{*}_{i2}+2c_{W}^{2}c_{2W}(V^{*}_{j1}V_{i1}U_{j2}U^{*}_{i2}+h.c.)],$
$\widetilde{\mathcal{C}}_{1\chi^{0}}^{ZZ}=$ $-2\frac{G^{2}}{4}(N^{*}_{j3}N_{i3}-N^{*}_{j4}N_{i4})^{2},$
$\widetilde{\mathcal{C}}_{2\chi^{0}}^{ZZ}=$ $2\frac{G^{2}}{4}(N^{*}_{j3}N_{i3}-N^{*}_{j4}N_{i4})^{2},$